how do I solve this equation?
8+a/2=4+2a

Answer by checkley77(12844)

8+a/2=4+2a
a/2-2a=4-8
(a-2*2a)/2=-4
(a-4a)/2=-4
a-4a=-4*2
-3a=-8
a=-8/-3
a=8/3 ans.
Proof:
8+(8/3)/2=4+2*8/3
8+8/3*1/2=4+16/3
8+8/6=4+16/3
(8*6+8)/6=(4*3+16)/3
(48+8)/6=(12+16)/3
56/6=28/3
6*28=56*3
168=168